# Formulation of entropy through work by Carnot machine and direct derivation of law of entropy non-decrease from Kelvin principle

### v2

@@ 3.2 Clausius entropy using isothermal heat
The outline of the above derivation is essentially the same as those in Secs. 2.2 and 2.3. An important difference is that the calculation of the work here is done by using the Helmholtz free energy difference for an isothermal process, while it is done by using the exergy for a temperature-changing process with a Carnot machine in Secs. 2.2 and 2.3. With this process using a Carnot machine, we are able to apply the Kelvin principle to the Clausius entropy Eq. (2) in terms of non-isothermal heat.
4 Conclusion
@@
We derived the principle of entropy non-decrease Eq. (1) from Kelvin principle Eq. (5) for both simple and compound systems without using Clausius inequality Eq. (3), where the process that utilizes the Carnot machine operating between the system and the heat reservoir played the essential role in the derivation. Our entropy Eq. (10) is equivalent to Clausius entropy Eq. (2) and Gyftopoulos-Beretta entropy Eq. (27), establishing a thermodynamic consistency. We hope that the present derivation elucidates the principle of thermodynamics. It is also of interest to apply the present formulation of entropy to determine adiabatic accessibility between nonequilibrium states [8], where Gyftopoulos-Beretta entropy serves as one expression to determine such accessibility.
Acknowledgments

### v5

@@ III.2 Clausius entropy using isothermal heat
The outline of the above derivation is essentially the same as those in Secs. II.2 and II.3. An important difference is that the calculation of the work here is done by using the Helmholtz free energy difference for a reversible isothermal process, while it is done by using the reversible work Eq. (11) for a reversible temperature-changing process with a Carnot machine in Secs. II.2 and II.3. With this process using a Carnot machine, we were able to apply the Kelvin-Planck principle to show the non-decrease of the entropy Eq. (10), which is essentially the same entropy as the Clausius entropy Eq. (2) in terms of non-isothermal heat.
III.3 Characterization of entropy as thermodynamic cost to create a system’s state from a reference state
Figure 3: A compound system at state $A_{1}$ and a simple system at state $A_{1^{\prime}}$ of the same internal energy and volume are created from a simple system at state $A_{0}$. An adiabatic process $A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{1^{\prime}}$, which occurs spontaneously after the removal of the insulated walls separating the subsystems of the compound system, is realized only when Eq. (34) is satisfied.
In light of the discussion so far, we may characterize entropy as the thermodynamic cost to create a system’s state from a reference state. To this end, let us consider a setup in Fig. 3: From a simple system at state $A_{0}$, a compound system at state $A_{1}$ constituted with $M$ subsystems and a simple system at state $A_{1^{\prime}}$ are created, where the internal energies and volumes of the states $A_{1}$ and $A_{1^{\prime}}$ are set to be the same. The thermodynamic cost, the work taken for creating each state by combining reversible adiabatic processes and reversible temperature-changing processes with a Carnot machine (Fig. 1), is calculated as the minus of the extracted work as
$-W_{0\to 1}^{\rm rev}=-(U_{0}-U_{1}-T_{o}(S_{0}-S_{1})),$ (32)
$-W_{0\to 1^{\prime}}^{\rm rev}=-(U_{0}-U_{1^{\prime}}-T_{o}(S_{0}-S_{1^{\prime}})),$ (33)
where $U_{1}=\sum_{j=1}^{M}U_{1,j}$ and $S_{1}=\sum_{j=1}^{M}S_{1,j}$, and $U_{1}=U_{1}^{\prime}$. Consider a cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{1^{\prime}}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$, where during $A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{1^{\prime}}$ the allowed adiabatic operation is to remove all the insulated walls separating the subsystems so that the compound system at state $A_{1}$ spontaneously changes to the simple system at state $A_{1^{\prime}}$; the work taken for this process is assumed to vanish as $-W_{1\to 1^{\prime}}^{\rm ad}=0$. The application of Kelvin-Planck principle to this cyclic process yields $W_{0\to 1}^{\rm rev}\leq W_{0\to 1^{\prime}}^{\rm rev}$, which can be expressed in terms of the work cost as
$-W_{0\to 1}^{\rm rev}\geq-W_{0\to 1^{\prime}}^{\rm rev}.$ (34)
Note that this is equivalent to the law of entropy non-decrease $S_{1}\leq S_{1^{\prime}}$ as confirmed from Eqs. (32) and (33) with $U_{1}=U_{1^{\prime}}$. The expression Eq. (34) implies that the spontaneous process from $A_{1}$ to $A_{1^{\prime}}$ occurs only when $A_{1}$ is more costly than $A_{1^{\prime}}$ to create from $A_{0}$ thermodynamically; the low-entropy (or ordered) structure is hard to create, which is easily seen by comparing Eqs. (32) and (33) with $U_{1}=U_{1^{\prime}}$. This may give a characterization of entropy as the thermodynamic cost to create one state from another, which makes it possible to distinguish between two macroscopic systems of the same internal energy and volume.
As an example, we consider entropy of an ideal gas. We create states $A_{1}(\{T_{1,1},V/2\},\{T_{1,2},V/2\})$ ($M=2$) and $A_{1^{\prime}}(T_{1^{\prime}},V)$ from a reference state $A_{0}(T_{0},V)$. The internal energy $U_{1}=U_{1,1}+U_{1,2}$ and $U_{1^{\prime}}$ coincide as no thermodynamic cost is assumed to be necessary to remove the insulated wall partitioning the two subsystems. From Eq. (10), the entropy change from $A_{0}$ to $A_{1}$ reads
[…]
where $C_{V}$ denotes the constant-volume heat capacity of the ideal gas. In the same manner, we obtain the entropy change from $A_{0}$ to $A_{1^{\prime}}$ as
[…]
where we used $T_{1^{\prime}}=(T_{1,1}+T_{1,2})/2$ obtained from $U_{1}=U_{1,1}+U_{1,2}=U_{1^{\prime}}$ in the last equality. From Eqs. (32) and (33), the works taken for these changes read
$-W_{0\to 1}^{\rm rev}=-(U_{0}-U_{1})-C_{V}T_{o}\ln\frac{\sqrt{T_{1,1}T_{1,2}}}{T_{0}},$ (37)
$-W_{0\to 1^{\prime}}^{\rm rev}=-(U_{0}-U_{1^{\prime}})-C_{V}T_{o}\ln\frac{T_{1,1}+T_{1,2}}{2T_{0}},$ (38)
respectively, where we note $U_{1}=U_{1^{\prime}}$. Because of $(x+y)/2\geq\sqrt{xy}$ for $x,y\geq 0$, we have $-W_{0\to 1}^{\rm rev}\geq-W_{0\to 1^{\prime}}^{\rm rev}$.
IV Conclusion
@@
We derived the law of entropy non-decrease Eq. (1) from the Kelvin-Planck principle Eq. (5) for both simple and compound systems without using the Clausius inequality Eq. (3), where the process that utilizes the Carnot machine operating between the system and the heat reservoir played the essential role in the derivation. Our entropy Eq. (10) is equivalent to the Clausius entropy Eq. (2) and the Gyftopoulos-Beretta entropy Eq. (27), establishing a thermodynamic consistency. We also characterized entropy as the thermodynamic cost to create a system’s state from a reference state. We hope that the present derivation elucidates the principle of thermodynamics. It is also of interest to apply the present formulation of entropy to nonequilibrium entropies that determine adiabatic accessibility between nonequilibrium states [8].
Appendix A Proof that Eq. (11) serves as the maximum work
We show that Eq. (11) serves as the maximum work extracted from a system changing from one equilibrium state $A_{0}$ to another equilibrium state $A_{1}$ with the aid of a heat reservoir at fixed temperature $T_{o}$. The argument is based on Sec. 3(b) of Ref. [8].
Consider a thermodynamic process where a system changes its state from one equilibrium state $A_{0}$ to another $A_{1}$. There is no need for the system’s temperature to match the fixed temperature $T_{o}$ of the heat reservoir during the process including the initial and final states $A_{0}$ and $A_{1}$. The entropy change of the total system, i.e., the system and the heat reservoir [13], reads
$S_{1}-S_{0}+\Delta S_{R}\geq 0,$ (39)
where $\Delta S_{R}$ denotes the entropy change of the heat reservoir. The minus of the internal energy change of the total system, which is equivalent to the work $W$ extracted from the system, reads
$W=U_{0}-U_{1}-\Delta U_{R}=U_{0}-U_{1}-T_{o}\Delta S_{R},$ (40)
where $\Delta U_{R}=T_{o}\Delta S_{R}$ denotes the internal energy change of the heat reservoir. By applying the inequality Eq. (39) to Eq. (40), we obtain
$W\leq U_{0}-U_{1}-T_{o}(S_{0}-S_{1}).$ (41)
In the case using a reversible isothermal process, the maximum work, the right-hand side of Eq. (41), is nothing but the Helmholtz free energy difference (sec. III.2). But this is not the only case, and as we see in Eq. (11), the equality is attained also for the reversible temperature-changing process with a Carnot machine.
Acknowledgements.